OpenStudy (donnie1999):

What is the value of f(x) = 9^x when x = -2?

OpenStudy (donnie1999):

@KaiserTheSlayer

OpenStudy (kaisertheslayer):

f(-2)=0.012345679012346

OpenStudy (jack1):

What is the value of f(x) = 9^x when x = -2? \(\Large f(x) = 9^x\) when x = -2? \(\Large f(-2) = 9^{-2}\)

OpenStudy (donnie1999):

the answer choices are a. 1/81 b. 81 c. 1/18 d. 18

OpenStudy (jack1):

\(\Large f(-2) = 9^{-2}\) \(\Large f(-2) = \frac 1 {9^{2}}\) \(\Large f(-2) = \frac 1 {81}\)

OpenStudy (kaisertheslayer):

a.

OpenStudy (kaisertheslayer):

need amymore?